# H-DES: a Quantum-Classical Hybrid Differential Equation Solver

### v1

@@ 4 Implementation and related complexity
In this section, we conduct a theoretical study of our algorithm to examine the complexity of each step. More precisely, we aim to use theoretical arguments to study the algorithm scaling properties (for both the quantum and classical parts) as a function of the inputs of the problem to be solved (e.g., the number of sample points, number of variables, and number of DEs). This analysis will allow us to ensure the stability and realism of our method when moving to more complex systems of DEs. In the following, we will mainly focus on establishing formulas to estimate the computational time.
@@ 4.1 Pseudo-code of the algorithm
7:$n_{s}$ : the number of sample points
8:$\varepsilon$ : the target precision for the error
9:$m_{iter}$ : the maximum number of iterations of the optimizer
@@ 4.2 Line-by-line analysis
  • In line 14, we compute the expectation value of the circuit for each observable. This computation is usually delegated to the quantum devices that decompose the observable and select the measurements to optimize this process. We discuss this in more detail in Appendix C.3, and this step takes a time $T_{expvalue}=n_{shots}\times(T_{VQC\_exec}+T_{measurement})+T_{post\_estim}$.
  • In line 15, we compute the loss function. This step is detailed in Appendix C.4 and takes a time $T_{loss}=|E|\times n_{s}\times(T_{square}+\sum_{f\in F}|H_{E}|T_{sum})$.
@@ 4.3 Scaling properties of the algorithm
An increase in the order of the DE mainly affects the observables. First, it can require the evaluation of additional derivatives of the Chebyshev polynomials (if we suppose that all lower orders of derivatives appear in the DEs) in the classical pre-processing step, but the increase is proportional to the number of samples $n_{s}$. In addition, it can require the evaluation of more expectation values but this only affects the post-processing of such expectation values at each iteration, because they can also be deduced from the expectation value of the Pauli observables that are combinations of $I$ and $Z$, and thus do not require additional circuit measurements. The same argument for expectation values can be used when we increase the number of terms in the DEs. However, in this latter case, the evaluation of the equations for computing the loss consequently increases (also linearly).
@@
When the depth of the circuit increases, more gates are executed, which linearly increases both the execution time and the number of parameters to optimize, quadratically affecting the performance of the optimizer. The idea behind a VQA is to work with shallow circuits and in our context of low-qubit circuits, choosing an appropriate Ansatz can ensure that the required depth for good expressivity is also low [DTYT22].
@@ 5 Results
In this section, we present the results obtained when applying our algorithm to solve different ODEs. We consider three examples: the first is a system of two coupled linear and first-order ODEs; the second is the damped harmonic oscillator, represented by a linear second-order ODE; and the third is a hypoelastic deformation problem modeled by two coupled non-linear and first-order ODEs.
We run our algorithm on the ideal QLM simulator ($qlm\_35$) provided by Eviden and give the size and depth of the circuit for each example. To generate a final state solution, we use the BFGS [NW06]*Sec. 6.1 optimizer to update the circuit angles $\theta$ within our hybrid loop. The following results are obtained by averaging over 100 random(uniform) initializations of the parameters of the circuits. We present the results in plots showing the solution function at each sample point as computed by both our solver and a classical (Mathematica DSolve function) solver. For each point, we also compute and plot the standard deviation. Finally, we present the validation scores associated with each solution, as defined in Sec. 2.6. We compute the validation score using a set $S_{v}$ of 100 points, linearly spaced in the interval $[0,0.95]$.
5.1 System of two coupled linear first-order ODEs
@@
5.2 Damped harmonic oscillator
@@
5.3 Material deformation application
The last example we consider to demonstrate the performance of our algorithm is a material deformation problem. We consider simple and coupled DEs to describe one-dimensional samples, such as a tensile test, where a thin strip is fixed to a grip and pulled at the right end. One of the goals of solving DEs for static material deformation is to compute the final displacement of the solid under a load. This involves determining the spatial difference for each point between the initial and deformed solids.
@@ 5.3 Material deformation application
[…]
@@ 6.1 Advantages of our approach
The first advantage of our method is that the number of qubits required is limited. First, the representativity power of our approach grows exponentially with the number of qubits, since the number of Chebyshev polynomials doubles when we add one qubit. With an already low number of qubits, one can represent a large range of continuous functions. Second, the qubit requirement grows linearly with the number of dimensions. The latter is a crucial point: since the number of variables is usually limited by spatial and temporal dimensions (3D$+$1D), the number of qubits needed to represent the solution function is also limited.
@@
In addition, as described before, good results using shallow circuits ($d=3$ for 1D MD) have been obtained. The expressivity provided by the parametrized quantum circuit allows us to find the solution without a need for deep circuits.
Finally, another advantage of our method is the lack of error in the computation of the derivatives. Since the same circuit is used to define the state representing the function and its derivatives, no approximation is needed to estimate them. For these reasons, we believe that the main expected advantage provided by the algorithm is its potential to return very precise solutions with a limited amount of quantum resources. It is also possible that this advantage translates into a reduction in energy consumption for a fixed precision when compared to classical solvers.
@@ 6.2 Differentiation from existing algorithms
One restricted type of equation (Burger, Vlasov, …) [GJFC, ESP19, OVKG22]
@@
(using FEM, FDM, QFT) [ZMSS21, WWL+, GRG22, CL20, FL18, OVKG22, LJM+20, CLO21, LJM+20]
@@
(Semi-)spectral approach [CL20]
Not NISQ friendly [Gai21, Ber14, CLO21, Bud21, Ste20, GRG22]
Other paradigm (annealing) [SS19, PSG+21]
@@
We can compare our work in particular to that of Kyriienko et al. [KPE21], who proposed a general ODE solver based on differentiable quantum circuits. In our approach, we encode the dependence on $x$ in the observables, while Kyriienko et al. encode it directly in the circuit using quantum feature maps. In their approach, the evaluation of the solution function is a non-linear process, allowing different choices of observables of quantum feature maps, but not based on a specific encoding or decomposition of the solution (i.e., it is not a spectral approach). For evaluating the solution function or its derivatives, we replace Chebyshev polynomials with their derivatives in the diagonal observables, while the approach of Kyriienko et al. involves reconstructing the derivatives by evaluating several circuits (changing the quantum feature map of the circuit used to evaluate the solution function), whose number grows exponentially with the order of the derivative. However, Kyriienko et al. do not discuss in detail how to solve systems of DEs, or partial DEs and the encoding of multi-dimensional variables, for which the cost and its scaling are also not precised.
@@ 7 Conclusion
The algorithm is designed to handle any linear or non-linear DE. As detailed in previous sections, its ability to be easily implemented in the case of partial and coupled DEs demonstrates the genericity of the approach and the versatility of our solver. Its scalability properties show that adding more complexity to the problem by introducing non-linear terms, more variables, or more functions to the system of DEs only linearly increases the number of resources (qubits or circuits) and parameters to optimize. In addition, with the expressivity of the solution function increasing exponentially with the number of qubits, and with shallow circuits needed to encode it, the algorithm is suitable for running on current quantum devices. We thus believe this algorithm constitutes an important step toward toward realistic applications of hybrid quantum algorithms
@@
The natural next steps will be to demonstrate the ability of our algorithm to perform in a noisy environment and to develop a full pipeline including error mitigation techniques. Another possible improvement of the algorithm will be to provide an initialization strategy for the circuit parameters, for instance by exploiting the expression or properties of the DEs to start the optimization process as close as possible to the actual solution. In the direction of having a generic solver, one can imagine that some parameters of the solver (the number of qubits, depth, number of iterations, number of sample points) can be automatically chosen to satisfy the desired precision by taking into account the type and form of the DE. Improvements to the algorithm can be made at different stages to optimize some aspects of pre and post processing, to more efficiently evaluate the expectation values of the diagonal observables, and to provide a loss function less prone to barren plateaus and local minima, but also by selecting an appropriate classical optimizer.
@@
In conclusion, the hybrid quantum algorithm proposed in this paper exhibits promising potential for practical application to various use cases despite the current limitations inherent to quantum computing. By combining classical computational methods with nascent quantum capabilities, this algorithm provides an original solution for solving generic partial DEs. In the NISQ era, where quantum computers face challenges related to stability and computational power, the significance of this hybrid approach relies on the potential span of diverse sectors, from material science to finance, where its deployment could help optimize processes such as material design and risk assessment. This hybrid algorithm not only serves as a bridge between theoretical quantum advancements and industrial applicability but also underscores the current need for accelerated integration of quantum technologies into industrial workflows. We believe that the adoption of such algorithms will signify a critical stride toward realizing the transformative potential of quantum computation in practical industrial contexts.
@@ A.1 Floating boundary handling
The main difference between this method and that including the boundary conditions directly within the loss function is that the first takes the boundary conditions into account exactly and with certitude, since the value of the solution function is manually fixed. The other solution can only perfectly ensure the boundary condition is met if the loss function is exactly equal to zero. Having a slight imperfection in the boundary condition fit can lead to very different solutions, especially in a chaotic setup. However, floating boundary conditions handling cannot be used when the boundary conditions are more complex.
@@ B.1 Modeling a given function
Let $f:\mathbb{R}\to\mathbb{R}$ be the function defined by $f(x)=2x-3$. Since $f$ is a linear polynomial in $x$ of degree 1, it can be exactly expressed using the two first Chebyshev polynomials, i.e.,
@@
We see clearly from this expression that $f$ can be modeled using the global scaling factor $\lambda_{f}$ and the two-qubit quantum state |$\ket{\psi_{f}}$ such that:
@@
|$\ket{\psi_{f}}$ $=\frac{1}{\sqrt{5}}\left(\sqrt{2}\ket{01}+\sqrt{3}\ket{10}\right).$ (43)
@@ B.3 Simple example in the multivariate case
Let |$\ket{\psi_{g}}\in(\mathbb{C}^{2})^{\otimes 6}$ be a six-qubit state and let $\lambda_{g}=7$ be a scalar used to model a bivariate solution function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, such that :
|$\ket{\psi_{g}}=\frac{1}{\sqrt{14}}\left(2\ket{001000}+2\ket{100010}+\sqrt{6}\ket{010111}\right)~{},$
@@
We choose to encode the $x$-dependence with three qubits and the $y$-dependence with two qubits. Therefore, we can divide any basis state |$\ket{i}$ of six-qubit systems as follows:
@@
with $\textrm{bin}(L_{1})=i_{1}i_{2}i_{3}$ and $\textrm{bin}(L_{2})=i_{4}i_{5}$. If we take for instance the first basis state appearing in |$\ket{\psi_{g}}$, we can associate it with the following multivariate Chebyshev polynomial after computation of the expectation value of the diagonal observable $O_{C}(x,y)$:
|$\underbrace{\ket{0}}_{+}\underbrace{\ket{010}}_{2}\underbrace{\ket{00}}_{0}~{}\xrightarrow{\langle O_{C}(x,y)\rangle}~{}\text{Cheb}(2,x)\text{Cheb}(0,y)=2x^{2}-1~{},$
@@ C.1 Generating the observables
11:         $\text{order}_{g}$leftarrow [ ]
@@
14:              for i from 0 to $12^{n-1}-$ 1 do
15:                  C.append(Cheb($i$, $x_{s}$, $g$))
@@
17:              observable $\leftarrow$ diag(C, --C)
@@
$O_{C}(x)=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x)\outerproduct{i}{i}\right)$
@@ C.2 Variational quantum circuit (VQC)
[…]
Figure 6: An example of our VQC with $n=5$ qubits and depth $d=2$.
There are two types of times attached to the VQC: the initialization of the circuit and the execution of the circuit. The initialization takes a time
$T_{VQC\_init}=(2n-1)\times d\times(T_{gate\_creation}+T_{instruction\_adding})$ (54)
and the execution of the circuit takes a time
$T_{VQC\_exec}=3\times d\times T_{gate\_exec}.$ (55)
One can remark that, even in the presence of several functions in the differential equations, all the circuits can be run in parallel. When the number of qubits and the depth are the same for all functions, we can also imagine only using a single circuit that is initialized once and just replacing the angles to evaluate one function or another.
C.3 Expectation values
In order to determine the value of the solution $f$ at a given point $x$, we compute the expectation value with respect to the diagonal observable $O_{C}(x)$ discussed above as follows:
$𝜓𝜃f_{\theta}(x)=\lambda~{}\expectationvalue{O_{C}(x)}{\psi(\theta)}.$ (56)
Though one has to deal only with diagonal matrices, evaluating the expectation values of several observables constitutes the costly step of the quantum part of the algorithm. The method of expressing the elements of the matrices is convenient since evaluating the expectation values of Pauli strings through a quantum computer is practical, especially if the Pauli strings commute, since in that case they can be measured simultaneously [KRP22]. The upper bound of the number of such strings per observable is $2^{n-1}$, with $n$ the number of qubits. So in the worst case scenario, in which no strings commute, one has $T_{eval}=2^{n-1}T_{measurement}$.
But in our case, our Pauli strings have a very specific form: they all start with a Pauli $Z$ gate and can then take all combinations of Pauli $I$ and $Z$. In that case, all the Pauli operators used to decompose our diagonal observables commute. Therefore, one can imagine many protocols and methods for efficiently evaluating the expectation values of such observables, even in the presence of noise [ZNCA, KSK+22, HKP21, CG23a, HBRM20, KIK+22, RB20, HKP20, KG22]. Once we know all the expectation values of the Pauli observables, we can reconstruct all expectation values of all diagonal observables (for all samples, for evaluating solution functions or their derivatives) in parallel with classical post-processing.
The evaluation of the expectation values consists of running the circuit and carrying out the right measurements for each shot, then post-processing them to reconstruct all desired expectation values, giving
$T_{expvalue}=n_{shots}\times(T_{VQC\_exec}+T_{measurement})+T_{post\_estim}.$ (57)
C.4 Loss function (error)
The value we want to minimize is:
where the notation is the same as that defined in Sec. 2.5.
We omit the term in the loss that concerns the boundary conditions and apply floating boundary conditions handling. Recall that we have:
  • $\theta$: the set of angles parametrizing our solution,
  • $E$: the set of differential equations,
  • $n_{s}$: the number of samples,
  • $F$: the set of functions,
  • $n_{f}$: the number of functions involved in the equations,
To compute this sum, we iterate over:
  1. 1.
    the set of differential equations $E$,
  2. 2.
    the set of samples ($n_{s}$ elements),
  3. 3.
    the set $H_{E}$ of all solution functions and their derivatives appearing in the differential equations.
Given this, the time to compute this loss value in the worst case (for all functions, with the function and its derivatives appearing in each equation) would be:
$T_{loss}=|E|\times n_{s}\times(T_{square}+\sum_{f\in F}|H_{E}|T_{sum}),$ (59)
where $T_{square}$ is the time required to square a floating point number and $T_{sum}$ is the time needed to sum two floating point numbers.
Appendix D Performance profile
Performance profiles, as described in [DM02], are a fairly common tool in optimization when it comes to comparing the performance of different algorithms, software, or methods for solving a problem. They allow for assessing both the speed and accuracy of an algorithm in relation to the proposed solutions. They can also be used to compare the performance of a given solver with different parameters and settings.
We denote the set of solvers to compare as $\mathcal{S}$ and the set of problems on which the solvers will be compared as $\mathcal{P}$. In the context of this paper, the problems to solve correspond to different sets of partial DEs. One then usually needs to choose a criterion to measure the performance or quality of the solver. It can be the number of evaluations of the solution function, the final value of the loss function, the precision of the result, or the number of iterations needed for a given precision. In our case, the validation score we defined previously is used as the measure of performance for a given solver.
We thus define the quantity $V_{p,sol}$, which is the validation score of the solver $sol\in\mathcal{S}$ for solving the problem $p\in\mathcal{P}$. In order to be able to compare several solvers, we need to have a common basis for comparison. As detailed in [DM02], for a given problem $p$, we compare the performance of each solver with the best performance over $\mathcal{S}$. We then define the performance ratio $r_{p,sol}$ as follows:
$r_{p,sol}=\frac{V_{p,sol}}{\min\limits_{sol\in\mathcal{S}}V_{p,sol}}~{}.$ (60)
Therefore, for a given problem $p$, we define the performance ratio as the validation score of the solver for this particular problem divided by the best validation score (the minimum) among all other solvers, still for the same problem. We determine that the solver could not solve the problem if the validation score is above the acceptance threshold. If a solver is not able to solve a problem, the performance ratio is set to a maximum value $r_{M}$.
We can now define the global performance of a solver $sol$ for all problems in $\mathcal{P}$ as the cumulative distribution function
$\rho_{sol}(\tau)=\frac{\text{card}\Big{(}\{p\in\mathcal{P}~{}|~{}r_{p,sol}\leq\tau\}\Big{)}}{\text{card}(\mathcal{P})}~{},$ (61)
where $\rho_{sol}(\tau)$ is the probability for the solver $sol$ that a performance ratio $r_{p,sol}$ is within a factor $\tau\in\mathbb{R}$ of the best possible ratio (which is 1). The function $\rho_{sol}$ can be seen as the (cumulative) distribution function for the performance ratio. To obtain the performance graph of the solver $sol$, we will represent the strictly increasing function $\rho_{sol}:\mathbb{R}\to[0,1]$ as a function of $\tau$.
The value $\rho_{sol}(1)$ is the probability that the solver $sol$ has a better validation score than all the other solvers. Therefore, if we are interested in precision, we can only compare the values of $\rho_{sol}(1)$ for all solvers in $\mathcal{S}$.
Furthermore, we recall that any performance ratio takes a value within $[1,r_{M}]$, and that $r_{p,sol}=r_{M}$ only if the problem $p$ was not solved by $sol$. We can deduce that $\rho_{sol}(r_{M})=1$ and that the quantity
$\rho_{sol}^{*}=\lim_{\tau\rightarrow{r_{M}}^{-}}\rho_{sol}(\tau)$ (62)
is the probability that $sol$ is able to solve a problem. Thus, if we are only interested in solvers that have the highest probability of success, we compare values of $\rho_{sol}^{*}$ for all solvers in $\mathcal{S}$ and keep that with the highest value. The value of $\rho_{sol}^{*}$ can be easily observed in a performance profile graph because the value of $\rho_{sol}$ stagnates for high values of $\tau$.
Figure 7: Example of a performance profile graph from [DM02]. The highest value of $\rho_{sol}(1)$ is here achieved by LOQO (0.61), while the highest value of $\rho_{sol}^{*}$ is apparently achieved by MINOS (0.8) for $\tau\in[1,10]$.

### v2

@@ 4 Implementation and related complexity
In this section, we provide preliminary elements for a theoretical study of our algorithm to examine the complexity of each step. More precisely, we aim to use theoretical arguments to study the algorithm scaling properties (for both the quantum and classical parts) as a function of the inputs of the problem to be solved (e.g., the number of sample points, number of variables, and number of DEs). The goal of this analysis is to provide initial insights on the stability and realism of our method when moving to more complex systems of DEs. In the following, we will mainly focus on the estimation of the computational time.
@@ 4.1 Pseudo-code of the algorithm
[…]
@@ 4.2 Line-by-line analysis
  • In line 14, we compute the expectation value of the circuit for each observable. This computation is usually delegated to the quantum devices that decompose the observable and select the measurements to optimize this process. We discuss this in more detail in Appendix LABEL:sub:expectation_values, and this step takes a time $T_{expvalue}=n_{shots}\times(T_{VQC\_exec}+T_{measurement})+T_{post\_estim}$.
  • In line 15, we compute the loss function. This step is detailed in Appendix LABEL:sub:error and takes a time $T_{loss}=|E|\times n_{s}\times(T_{square}+\sum_{f\in F}|H_{E}|T_{sum})$.
@@ 4.3 Scaling properties of the algorithm
An increase in the order of the DE mainly affects the observables. First, it can require the evaluation of additional derivatives of the Chebyshev polynomials (if we suppose that all lower orders of derivatives appear in the DEs) in the classical pre-processing step, but the increase is proportional to the number of samples $n_{s}$. In addition, it can require the evaluation of more expectation values, but this only affects the post-processing of such expectation values at each iteration, because they can also be deduced from the expectation value of the Pauli observables that are combinations of $I$ and $Z$, and thus do not require additional circuit measurements. The same argument for expectation values can be used when we increase the number of terms in the DEs. However, in this latter case, the evaluation of the equations for computing the loss consequently increases (also linearly).
@@
When the depth of the circuit increases, more gates are executed, which linearly increases both the execution time and the number of parameters to optimize, quadratically affecting the performance of the optimizer. The idea behind a VQA is to work with shallow circuits and, in our context of low-qubit circuits, choosing an appropriate Ansatz can ensure that the required depth for good expressivity is also low [DTYT22].
@@ 5 Results
In this section, we present the results obtained when applying our algorithm to solve different ODEs. We first consider three examples running on emulator: the first is a system of two coupled linear and first-order ODEs; the second is the damped harmonic oscillator, represented by a linear second-order ODE; and the third is a hypoelastic deformation problem modeled by two coupled non-linear and first-order ODEs. Finally we execute our algorithm on a real quantum backend to solve a linear first order ODE.
5.1 Execution on a quantum emulator
We run our algorithm on the ideal QLM simulator (qlm_35qlm\_35) provided by Eviden and give the size and depth of the circuit for each example. To generate a final state solution, we use the BFGS [NW06]*Sec. 6.1 optimizer to update the circuit angles $\theta$ within our hybrid loop. The following results are obtained by averaging over 100 random(uniform) initializations of the parameters of the circuits. We present the results in plots showing the solution function at each sample point as computed by both our solver and a classical (Mathematica DSolve function) solver. For each point, we also compute and plot the standard deviation. Finally, we present the validation scores associated with each solution, as defined in Sec. 2.6. We compute the validation score using a set $S_{v}$ of 100 points, linearly spaced in the interval $[0,0.95]$.
5.1.1 System of two coupled linear first-order ODEs
@@
5.1.2 Damped harmonic oscillator
@@
5.1.3 Material deformation application
The next example we consider is a simple material deformation problem. We consider simple and coupled DEs to describe one-dimensional samples, such as a tensile test, where a thin strip is fixed to a grip and pulled at the right end. One of the goals of solving DEs for static material deformation is to compute the final displacement of the solid under a load. This involves determining the spatial difference for each point between the initial and deformed solids.
@@ 5.1.3 Material deformation application
$\left\{\begin{aligned} ~\frac{du}{dx}=\epsilon_{xx}(\sigma_{xx})\\ ~\frac{d\sigma_{xx}}{dx}+b_{x}=0\end{aligned}\right$.~, (34)
@@
5.2 Running on real quantum hardware
The last example we consider to demonstrate the performance of our algorithm is the resolution of a simple linear first-order ordinary differential equation on a real quantum processing unit. The goal is to showcase the capabilities of the algorithm when executed on today’s quantum computers on a simple enough equation. Many quantum approaches proposed in the literature for solving ODEs and PDEs fail to provide satisfying results when executed on real quantum hardware, even for relatively simple examples. In our case, we consider a canonical example described by the following ODE:
$f^{\prime}(x)-f(x)=0~,$ (36)
with the boundary condition $f(0)=1$, on the interval $[0,0.91]$. The resolution was performed using 16 sample points uniformly distributed on the interval. We run the algorithm using 3 qubits, with a typical $R_{y}$ with $CNOT$ (in cascade) ansatz of depth 2, and 20 iterations of the SLSQP optimizer. At each iteration, we evaluated expectation values with respect to the generated state using 20000 shots, with an analytical computation of the gradient. The boundary condition was imposed exactly using floating boundary handling (see Appendix A.1). The execution was made on IBM Quantum backend Brisbane (January 2025), with Eagle r3 processor type, on which the circuit was transpiled using Qiskit 1.2.4.
In Fig. 6, we compare the attempted solution with the analytical solution. We obtain a global validation score of $V=(1.57\cdot 10^{-2},1.46\cdot 10^{-4})$ and a final loss equal to $0.052$. This constitutes the first successful resolution of a differential equation on a NISQ device using the H-DES algorithm.
Figure 6: The exact (blue solid) versus attempted solutions (orange dashed) for $f(x)$ after 20 iterations of the SLSPQ optimizer, using analytic computation of the gradient. The fit was obtained by minimizing the loss function over 16 points equally spaced in the $x=[0,0.91]$ domain. The area at the bottom shows the absolute error (between the attempted and target solutions) associated with each sample point.
@@ 6.1 Advantages of our approach
The first advantage of our method is that the number of qubits required is limited. First, the representativity power of our approach grows exponentially with the number of qubits, since the number of Chebyshev polynomials doubles when we add one qubit. With an already low number of qubits, one can represent a large range of continuous functions. Second, the qubit requirement grows linearly with the number of dimensions. The latter is a crucial point: since the number of variables is usually limited by spatial and temporal dimensions (3D++1D), the number of qubits needed to represent the solution function is also limited.
@@
In addition, as described before, good results using shallow circuits ($d=3$ for 1D MD) have been obtained. The expressivity provided by the parametrized quantum circuit allows us to find the solution without a need for deep circuits. Preliminary empirical observations on various equations seem to indicate that the depth is scaling at most linearly (sometimes logarithmically) with the number of qubits.
Another advantage of our method is the lack of error in the computation of the derivatives. Since the same circuit is used to define the state representing the function and its derivatives, no approximation is needed to estimate them. For these reasons, we believe that the main expected advantage provided by the algorithm is its potential to return very precise solutions with a limited amount of quantum resources. It is also possible that this advantage translates into a reduction in energy consumption for a fixed precision when compared to classical solvers.
Finally, a general advantage is the flexibility allowed by the H-DES approach, providing a general framework for PDEs resolution, where the ansatz, the initial state, the choice of the loss function, the choice of the observable, the distribution of the sample (collocation) points (for optimization and for reconstruction of the solution) and the choice of the classical optimizer are customizable.
@@ 6.2 Differentiation from existing algorithms
One restricted type of equation (Burger, Vlasov, …) [GJFC, ESP19, OVKG22, LiYin2025, quanscient_press_release]
@@
Solves only first order PDEs [Pool2024]
@@
(using FEM, FDM, QFT, …) [ZMSS21, WWL+, GRG22, CL20, FL18, OVKG22, LJM+20, CLO21, LJM+20]
@@
(Semi) Spectral approach [CL20, Schillo2025]
Not NISQ friendly [Gai21, Ber14, CLO21, Bud21, Ste20, GRG22, SONG2025, LiYin2025, Alipanah2025]
Other paradigm (annealing,…) [SS19, PSG+21]
@@
We can highlight the main differences between our work and what was proposed by Kyriienko et al. in [KPE21], who proposed a general ODE solver based on differentiable quantum circuits. In our approach, we encode the dependence on $x$ in the observables, while Kyriienko et al. encode it directly in the circuit using quantum feature maps. In their approach, the evaluation of the solution function is a non-linear process, allowing different choices of observables or quantum feature maps, but not based on a specific encoding or decomposition of the solution (i.e., it is not a spectral decomposition). For evaluating the solution function or its derivatives, we replace Chebyshev polynomials (or any other orthogonal basis) with their derivatives in the observables, while the approach of Kyriienko et al. involves reconstructing the derivatives by evaluating several circuits (changing the quantum feature map of the circuit used to evaluate the solution function), whose number grows exponentially with the order of the derivative. However, Kyriienko et al. do not discuss in detail how to solve systems of DEs, or partial DEs and the encoding of multi-dimensional variables, for which the cost and its scaling are also not precised.
@@ 7 Conclusion
The algorithm is designed to handle any linear or non-linear DE, and is thought to solve various use cases and multiphysics simulations. As detailed in previous sections, its ability to be easily implemented in the case of partial and coupled DEs demonstrates the genericity of the approach and the versatility of our solver. Its scalability properties show that adding more complexity to the problem by introducing non-linear terms, more variables, or more functions to the system of DEs only linearly increases the number of resources (qubits or circuits) and parameters to optimize. In addition, with the expressivity of the solution function increasing exponentially with the number of qubits, and with shallow circuits needed to encode it, the algorithm is suitable for running on current quantum devices. We thus believe this algorithm constitutes an important step toward toward realistic applications of hybrid quantum algorithms
@@
The natural next steps will be to demonstrate the ability of our algorithm to perform on more complicated simulations and models, moving towards highly dimensional, stiff, chaotic, turbulent, discontinuous and stochastic phenomena. Another possible improvement of the algorithm will be to provide an initialization strategy for the circuit parameters, for instance by exploiting the expression or properties of the DEs to start the optimization process as close as possible to the actual solution. In the direction of having a generic solver, one can imagine that some parameters of the solver (the number of qubits, depth, number of iterations, number of sample points) can be automatically chosen to satisfy the desired precision by taking into account the type and form of the DE. Improvements to the algorithm can be made at different stages to optimize some aspects of pre and post processing, to more efficiently evaluate the expectation values of the observables, and to provide a loss function less prone to barren plateaus and local minima, but also by selecting an appropriate classical optimizer.
@@
In conclusion, the hybrid quantum algorithm proposed in this paper exhibits promising potential for practical application to various use cases despite the current limitations inherent to quantum computing. By combining classical computational methods with nascent quantum capabilities, this algorithm provides an original solution for solving generic partial DEs. In the NISQ era, where quantum computers face challenges related to stability and computational power, the significance of this hybrid approach relies on the potential span of diverse sectors, from material science to finance, including fluid dynamics and semiconductors, where its deployment could help optimize processes such as industrial design and risk assessment. This hybrid algorithm not only serves as a bridge between theoretical quantum advancements and industrial applicability but also underscores the current need for accelerated integration of quantum technologies into industrial workflows. We believe that the adoption of such algorithms will signify a critical stride toward realizing the transformative potential of quantum computation in practical industrial contexts.
@@ A.1 Floating boundary handling
The main difference between this method and that including the boundary conditions directly within the loss function is that the first takes the boundary conditions into account exactly and with certitude, since the value of the solution function is manually fixed. The other solution can only perfectly ensure the boundary condition is met if the loss function is exactly equal to zero. Having a slight imperfection in the boundary condition fit can lead to very different solutions, especially in a chaotic setup. However, floating boundary conditions handling cannot be used when the boundary conditions are more complex, and a more general interpolation scheme would be needed.
@@ B.1 Modeling a given function
Let $f:\mathbb{R}\to\mathbb{R}$ be the function defined by f(x)=2x3f(x)=2x-3. Since $f$ is a linear polynomial in $x$ of degree 1, it can be exactly expressed using the two first Chebyshev polynomials, i.e.,
@@
[…]
@@
[…]
@@ B.3 Simple example in the multivariate case
[…]
@@
[…]
@@ C.1 Generating the observables
11:   $\text{order}_{g}$ $\leftarrow$ [ ]
@@
14:     for i from 0 to $2^{n-1}-1$ do
15:      C.append(Cheb(ii, $x_{s}$, $g$))
@@
17:     observable $\leftarrow$ diag(C, -C)
@@
[…]